# arXiv:2603.02132v1[hep-ph] 02 Mar 2026

### v1

@@ Abstract
Sub-GeV thermal dark matter weakly interacting with the Standard Model through vector-portal mediators provides a well-motivated and predictive framework that remains challenging to probe with conventional direct detection experiments. Motivated by the rapid development of neutrino facilities based on spallation neutron sources, we study the sensitivity of future coherent elastic neutrino–nucleus scattering experiments to light dark matter produced in neutral pion decays. We consider scalar dark matter interactions mediated by two different vector portals, a generic dark photon and a baryophilic vector mediator. The neutral pion yield is calculated through a GEANT4 simulation and the results are compared with those obtained with the Sandford-Wang parametrization. We show that predictions based on either approach do not produce significant differences. Our results demonstrate that upcoming low-threshold neutrino detectors at the European Spallation Source (ESS), the Japan Proton Accelerator Research Complex (J-PARC) and the China Spallation Neutron Source (CSNS) will test regions in parameter space not yet explored. We point out that these facilities will strengthen the global experimental program searching for secluded sectors.
@@ I Introduction
To estimate the DM production rate, the neutral pion flux can be determined either through a dedicated simulation with existing toolkits like GEANT4 Agostinelli and others (2003) or, at the proton energies relevant in these installations, through the semi-empirical Sandford and Wang parametrization Sanford and Wang (1967). We compare the results from both approaches and then adopt the GEANT4-based fluxes for the full numerical analysis. Our simulations are validated against publicly available results from the COHERENT Collaboration at the Spallation Neutron Source (SNS) Akimov and others (2022a). For the ESS and J-PARC facilities, we consider medium-scale germanium, gaseous xenon, and CsI detectors Baxter and others (2020); Collar and others (2025), while for CSNS we assume a large-scale CsI detector Xu and others (2009). Under realistic assumptions on detector performance and backgrounds, we show that these setups can test previously unexplored regions of parameter space, including areas that remain challenging for other existing or proposed experimental searches.
@@ II.1 Model 1: Generic dark photon scenario
More generally, viable scenarios are those in which late-time annihilation is suppressed. This condition is naturally fulfilled when $p$-wave processes control freeze-out, as in the scalar DM case. By contrast, in the simplest fermionic DM realization with vector couplings, the relic density is typically set by $s$-wave $s$-channel annihilation, which remains efficient at late times and is therefore severely constrained by CMB data. Nonetheless, consistent fermionic constructions also exist. For example, pseudo-Dirac DM models, where the relic abundance is governed by co-annihilation between nearly degenerate states, avoid these bounds because the heavier state becomes Boltzmann suppressed at late times, and so the co-annihilation process Carrillo González and Toro (2022); Brahma et al. (2024); Garcia et al. (2025). Majorana fermion DM provides another viable option: its axial coupling to the vector mediator leads again to $p$-wave–suppressed annihilation Berlin et al. (2019). Finally, asymmetric fermionic DM scenarios can also evade recombination limits, since an early depletion of the anti-DM population renders late-time annihilation and the associated energy injection ineffective Berlin et al. (2019).
@@ III Production and detection of DM: Sanford and Wang parametrization
The positively charged pions decay mainly at rest, producing low-energy neutrinos, while negatively charged pions are absorbed by nuclei via Coulomb interactions (almost entirely at 1 GeV). Hence, their contribution to the neutrino flux is subdominant, although increasing with proton beam energy Akimov and others (2022a). On the other hand, neutral pions decay almost exclusively into two photons, which are absorbed within the target material given their short mean free path in highly dense environments [^3] For both mercury and tungsten the mean free path for a 50 MeV photons is about 1 cm 84.. The mechanism by which photons are absorbed is mainly pair creation in the electromagnetic field of nuclei, $\gamma+Z\to e^{+}+e^{-}+Z$. Most photons create pairs that are then further reprocessed by the target material either through electron-positron pair annihilation or atomic electron capture. Hence, electromagnetic activity is strongly suppressed in a thick target. This picture, however, is modified in the presence of a dark photon. If $\pi^{0}\to\gamma A^{\prime}$ is allowed through kinetic mixing, the produced dark photon can then decay into DM pairs. Since the dark photon and the DM feebly interact with matter, no process can induce their absorption. Then, the DM particle can escape the target material, and could be detected through nuclear recoil signals as those induced by neutrinos through CEνNS.
@@ III.1 Neutral pion PDF’s: Analytical versus simulation-based results
Note that, when sampling from Eq. (9), the material target plays no role, only the proton beam energy and the range over which $T_{\pi^{0}}$ run are relevant. Also, the S & W parametrization is well defined for $0\leq\theta_{\pi^{0}}\leq\pi/2$. For values above $\pi/2$ we adopt the following convention 91
@@ V Results
In this section, we present the projected sensitivities for sub-GeV DM for the experimental configurations discussed above. Figure 8 shows the 90% CL exclusion regions in the $(m_{A^{\prime}},Y)$ parameter space for detectors deployed at the ESS and J-PARC with CsI (solid contours), xenon (dotted) and germanium (dashed contours) targets, as well as at the CSNS with a CsI target (green solid contour). To compare with results in the literature, we stick to the following values: $m_{A^{\prime}}=3\,m_{\phi}$ and $\alpha_{D}=0.5$  [^9] Note that theoretical considerations on the running of the dark fine structure constant imply $\alpha_{D}\lesssim 0.1-0.5$ Davoudiasl and Marciano (2015).. The left panels in Fig. 8 show projected sensitivities for detectors placed on-axis, where the DM angular distribution peaks. The right panels, instead, display expected sensitivities for detectors located $90^{\circ}$ off-axis with respect to the proton beam direction (and inline with the location of the CsI detector in the COHERENT experiment). In all scenarios, we assume an exposure time of 3 years, corresponding to 208 days (3600 hours) per calendar year at the ESS and CSNS (J-PARC). In the analysis of the J-PARC and CSNS facilities, we have included a cut on the arrival time of neutrino events to reduce the CEνNS and SSB backgrounds as discussed in Sec. IV.1. We refer the reader to App. A where we discuss how our sensitivities are expected to change by varying some of the assumptions on backgrounds’ uncertainties in the statistical analysis, and what the impact of not including the timing information on the neutrino events is.
@@
For comparison, we display dark-pink contours corresponding to existing bounds from CEνNS-based experiments, namely the COHERENT CsI detector Akimov and others (2023) (solid) and CCM Aguilar-Arevalo and others (2022a) (dashed). We also include additional constraints (light-gray shaded region) obtained from MiniBooNE Aguilar-Arevalo and others (2018), NA64 Andreev and others (2023), LSND Auerbach and others (2001); Batell et al. (2009), E137 Batell et al. (2014b), and BaBar Lees and others (2017). These limits, however, should be interpreted with care, as they probe different processes and interaction channels. In particular, they are not directly related to CEνNS. MiniBooNE and LSND, although neutrino-beam experiments, are based on quasi-elastic scattering and inverse $\beta$ decay, respectively, while E137 and BaBar rely on $e^{-}-\text{Al}$ fixed-target and $e^{+}e^{-}$ collider searches. Consequently, the comparison is indicative rather than strictly equivalent. We further show the preferred curves in the $Y$ parameter corresponding to the observed relic DM abundance for scalar, pseudo-Dirac, and Majorana dark matter Berlin et al. (2019). As discussed above, these lines represent well-motivated target sensitivities: below them, the DM annihilation cross section is insufficient to reduce the relic abundance to the observed value. Above them, the DM would be underabundant and require an additional candidate to explain the total observed relic density.
@@
As can be seen, future spallation-based neutrino experiments have the potential to significantly improve existing constraints, particularly in the mediator mass range $8~\mathrm{MeV}\lesssim m_{A^{\prime}}\lesssim 100~\mathrm{MeV}$. In this region, the projected sensitivities can surpass the thermal relic density target for a scalar DM candidate by up to an order of magnitude, with optimal performance around $m_{A^{\prime}}\sim\mathcal{O}(10~\mathrm{MeV})$ for the J-PARC and ESS Ge detectors. Indeed, at the ESS and J-PARC, germanium-based detectors provide the strongest constraints, especially in the low-mass regime, reaching sensitivities down to $Y\sim 2\times 10^{-12}$. In the case of the CsI detector at these facilities, it is clear that doubling its volume (as expected at J-PARC Collar and others (2025)), leads to a non-negligible improvement in sensitivity reach, as visible by comparing the J-PARC vs the ESS panels in Fig. 8. Despite its much larger mass, the CsI detector at the CSNS is slightly less constraining due to its higher operation threshold together with the higher SSB background, but still it allows to probe $Y\sim 2\times 10^{-11}$ at tens of MeV dark photon masses, beyond the relic density benchmark. In the low-mass regime, sensitivities tend to saturate mostly due to detector limitations, namely the loss of efficiency at low recoil energies.
@@
On the other side of the parameter space instead, the sensitivity of spallation neutron source experiments rapidly reduces for mediator masses above $m_{A^{\prime}}\sim 100~\mathrm{MeV}$, reflecting the kinematic constraint for DM production through neutral pion decays. Access to heavier mediator masses require higher energy proton beams capable of producing heavier neutral mesons. Depending on the proton beam collision energy, heavier mesons can be produced, thereby extending the experimental reach to larger values of $m_{A^{\prime}}$. In addition, DM production via proton Bremsstrahlung could further expand the accessible parameter space deNiverville et al. (2017); Blümlein and Brunner (2014). Extending up to the energy regime of SHIP and FASER, the Drell-Yang production channel is also available. Spallation neutron source facilities provide a unique production channel that facilitates the “deconstruction” of the physics underlying a DM signal. High statistics combined with high background discrimination power, enabled by timing spectra, make these facilities a unique environment for light secluded sectors. Altogether, these considerations highlight the strengths of spallation neutron source facilities in probing light dark sectors. Information stemming from these measurements are therefore complementary to those at high-energy accelerator facilities.
@@
Next, in Fig. 9 we compare our projected exclusion reach in the $(m_{A^{\prime}},Y)$ plane for the dark photon portal scenario with the sensitivities expected at other future facilities. Our projections assume on-axis detector configurations and three years of data taking. We include sensitivity estimates for two proposed upgrades of COHERENT optimized for CEνNS detection: a 700 kg undoped cryogenic CsI detector Barbeau and others (2024) (pink, dashed) and a 10-tonne LAr detector Barbeau and others (2024) (pink, dotted). We also show projections for the first phase of LDMX Åkesson and others (2018); Batell and others (2022) (light blue, dashed), the movable near detector of DUNE-PRISM with a 75-tonne LAr target De Romeri et al. (2019) (light green, dashed), and a five-year run of SHiP Ahdida and others (2021); Batell and others (2022) (coral, dashed). These comparisons should also be interpreted with care. LDMX and SHiP are dedicated DM experiments, while DUNE operates at higher beam energies and primarily probes $\text{DM}-e^{-}$ scattering. Moreover, the projected sensitivities of the future COHERENT upgrades benefit from substantially larger detector masses. Despite this, our results show that experiments at ESS, J-PARC, and CSNS can competitively explore a broad region of parameter space beyond current limits. In particular, they are sensitive to mediator masses in the range $m_{A^{\prime}}\simeq 20$9090 MeV, reaching nearly an order of magnitude below the thermal relic target for scalar DM.
@@
All the sensitivities presented up to this point in this section have been derived from the DM event rate computed using our GEANT4-based simulations, as described in Sec. IV. As anticipated, for the ESS-Xe detector we repeated the analysis employing instead the neutral pion PDF obtained from the S & W parametrization. For dark photon masses $m_{A^{\prime}}\sim 10$ MeV, the resulting DM energy distributions can differ by up to a factor of two, whereas for $m_{A^{\prime}}\sim 100$ MeV the discrepancy decreases to below a factor of 1.3. Despite these normalization differences, the spectral shapes are remarkably consistent between the two approaches. Overall, the S & W parametrization systematically predicts fewer events than GEANT4, leading to slightly more conservative sensitivity projections. These differences propagate to the final exclusion contours, implying that our results are robust at the level of $\mathcal{O}(50\%)$ when comparing the full simulation with the analytical approximation. It is worth noting that uncertainties of comparable size may arise from the simulation framework itself, for instance, when switching between hadronic models in GEANT4. We present the result of this comparison in Fig. 10 where we compare the projected sensitivities for the Xe detector at the ESS obtained using the two approaches discussed in this work: the neutral pion PDF derived from the S & W parametrization and that extracted from our dedicated GEANT4 simulations. As previously noted, the S & W parametrization systematically predicts a lower DM yield than GEANT4. This translates into correspondingly weaker constraints, with sensitivity projections that are typically up to a factor of $\sim 2$ more conservative across the relevant mediator mass range.
@@ VI Conclusions
We have considered representative detector configurations, including planned gaseous xenon (20 kg), high-purity germanium (7 kg), and cryogenic CsI scintillating detectors (22.5 kg and 44 kg) at ESS and J-PARC, as well as a 300 kg CsI detector under construction at CSNS. A detailed discussion of the neutral pion flux calculation at these facilities has been presented. We compared two approaches: a dedicated GEANT4 simulation of proton collisions on mercury and tungsten targets, and the Sanford-Wang parametrization of charged-pion differential cross sections. We find that, despite minor spectral differences, both methods yield consistent results within the intrinsic uncertainties of the overall calculation. Our analysis focuses on a sub-GeV scalar dark matter candidate in the minimal dark photon portal framework, as well as on a baryophilic scenario based on gauged baryon number. Since the considered detectors are optimized for nuclear recoil measurements, we have studied coherent elastic dark matter-nucleus scattering signals.
Projected sensitivities have been presented in the $m_{A^{\prime}}$-$Y$ plane. Our results show that spallation-source neutrino experiments can probe a substantial region of previously unexplored parameter space, particularly in the mediator mass range $8~\mathrm{MeV}\lesssim m_{A^{\prime}}\lesssim 100~\mathrm{MeV}$, extending well beyond the thermal relic target for scalar dark matter. Even stronger sensitivities are obtained in baryophilic mediator scenarios (gauged baryon number models), as detailed in App. A. Our results highlight the strong complementarity between spallation-source facilities, optimized for CEνNS measurements, and other accelerator-based dark matter searches. By leveraging their nuclear recoil capabilities, these experiments offer a uniquely powerful environment for dark matter searches in the MeV mass range.

### v2

# Sensitivity to sub-GeV dark matter in forthcoming spallation-source neutrino experiments

@@ Abstract
Sub-GeV thermal dark matter weakly interacting with the Standard Model through vector-portal mediators provides a well-motivated and predictive framework that remains challenging to probe with conventional direct detection experiments. Motivated by the rapid development of neutrino facilities based on spallation neutron sources, we study the sensitivity of future coherent elastic neutrino–nucleus scattering experiments to light dark matter produced in neutral pion decays. We consider scalar dark matter interactions mediated by two different vector portals, a generic dark photon and a baryophilic vector mediator. The neutral pion yield is calculated through a GEANT4 simulation and the results are compared with those obtained with the Sandford-Wang parametrization. We show that predictions based on either approach do not produce significant differences. Our results demonstrate that upcoming low-threshold neutrino detectors at the European Spallation Source (ESS), the Japan Proton Accelerator Research Complex (J-PARC) and the China Spallation Neutron Source (CSNS) may test regions in parameter space not yet explored, or be competitive with existing bounds. We point out that these facilities will strengthen the global experimental program searching for secluded sectors.
@@ I Introduction
To estimate the DM production rate, the neutral pion flux can be determined either through a dedicated simulation with existing toolkits like GEANT4 Agostinelli and others (2003) or, at the proton energies relevant in these installations, through the semi-empirical Sandford and Wang parametrization Sanford and Wang (1967). We compare the results from both approaches and then adopt the GEANT4-based fluxes for the full numerical analysis. Our simulations are validated against publicly available results from the COHERENT Collaboration at the Spallation Neutron Source (SNS) Akimov and others (2022a). For the ESS and J-PARC facilities, we consider medium-scale germanium, gaseous xenon, and CsI detectors Baxter and others (2020); Collar and others (2025), while for CSNS we assume a large-scale CsI detector Xu and others (2009). Under realistic assumptions on detector performance and backgrounds, we show that these setups can be competitive with other experiments in testing regions of parameter space close or beyond the relic-density target, including areas that remain challenging for other existing or proposed experimental searches.
@@ II.1 Model 1: Generic dark photon scenario
More generally, viable scenarios are those in which late-time annihilation is suppressed. This condition is naturally fulfilled when $p$-wave processes control freeze-out, as in the scalar DM case. By contrast, in the simplest fermionic DM realization with vector couplings, the relic density is typically set by $s$-wave $s$-channel annihilation, which remains efficient at late times and is therefore severely constrained by CMB data. Nonetheless, consistent fermionic constructions also exist. For example, pseudo-Dirac DM models, where the relic abundance is governed by co-annihilation between nearly degenerate states, avoid these bounds because the heavier state becomes Boltzmann suppressed at late times, and so the co-annihilation process Carrillo González and Toro (2022); Brahma et al. (2024); Garcia et al. (2025). Majorana fermion DM provides another viable option: its axial coupling to the vector mediator leads again to $p$-wave–suppressed annihilation Berlin et al. (2019). Finally, asymmetric fermionic DM scenarios can also evade recombination limits, since an early depletion of the anti-DM population renders late-time annihilation and the associated energy injection ineffective Berlin et al. (2019). These realizations seem also to be more favorable in light of recent direct-detection data, see, e.g., Krnjaic (2025); Cheek et al. (2025); Han and Krnjaic (2026).
@@ III Production and detection of DM: Sanford and Wang parametrization
The positively charged pions decay mainly at rest, producing low-energy neutrinos, while negatively charged pions are absorbed by nuclei via Coulomb interactions (almost entirely at 1 GeV). Hence, their contribution to the neutrino flux is subdominant, although increasing with proton beam energy Akimov and others (2022a). On the other hand, neutral pions decay almost exclusively into two photons, which are absorbed within the target material given their short mean free path in highly dense environments [^3] For both mercury and tungsten the mean free path for a 50 MeV photons is about 1 cm 88.. The mechanism by which photons are absorbed is mainly pair creation in the electromagnetic field of nuclei, $\gamma+Z\to e^{+}+e^{-}+Z$. Most photons create pairs that are then further reprocessed by the target material either through electron-positron pair annihilation or atomic electron capture. Hence, electromagnetic activity is strongly suppressed in a thick target. This picture, however, is modified in the presence of a dark photon. If $\pi^{0}\to\gamma A^{\prime}$ is allowed through kinetic mixing, the produced dark photon can then decay into DM pairs. Since the dark photon and the DM feebly interact with matter, no process can induce their absorption. Then, the DM particle can escape the target material, and could be detected through nuclear recoil signals as those induced by neutrinos through CEνNS.
@@ III.1 Neutral pion PDF’s: Analytical versus simulation-based results
Note that, when sampling from Eq. (9), the material target plays no role, only the proton beam energy and the range over which $T_{\pi^{0}}$ run are relevant. Also, the S & W parametrization is well defined for $0\leq\theta_{\pi^{0}}\leq\pi/2$. For values above $\pi/2$ we adopt the following convention 95
@@ V Results
In this section, we present the projected sensitivities for sub-GeV DM for the experimental configurations discussed above. Figure 8 shows the 90% CL exclusion regions in the $(m_{\phi},Y)$ parameter space for detectors deployed at the ESS and J-PARC with CsI (solid contours), xenon (dotted) and germanium (dashed contours) targets, as well as at the CSNS with a CsI target (green solid contour). To compare with results in the literature, we stick to the following values: $m_{A^{\prime}}=3\,m_{\phi}$ and $\alpha_{D}=0.5$  [^9] Note that theoretical considerations on the running of the dark fine structure constant imply $\alpha_{D}\lesssim 0.1-0.5$ Davoudiasl and Marciano (2015).. The left panels in Fig. 8 show projected sensitivities for detectors placed on-axis, where the DM angular distribution peaks. The right panels, instead, display expected sensitivities for detectors located $90^{\circ}$ off-axis with respect to the proton beam direction (and inline with the location of the CsI detector in the COHERENT experiment). In all scenarios, we assume an exposure time of 3 years, corresponding to 208 days (3600 hours) per calendar year at the ESS and CSNS (J-PARC). In the analysis of the J-PARC and CSNS facilities, we have included a cut on the arrival time of neutrino events to reduce the CEνNS and SSB backgrounds as discussed in Sec. IV.1. We refer the reader to App. A where we discuss how our sensitivities are expected to change by varying some of the assumptions on backgrounds’ uncertainties in the statistical analysis, and what the impact of not including the timing information on the neutrino events is.
@@
For comparison, we display dark-pink contours corresponding to existing bounds from CEνNS-based experiments, namely the COHERENT CsI detector Akimov and others (2023) (solid) and CCM Aguilar-Arevalo and others (2022a) (dashed). We also include additional constraints (light-gray shaded region) obtained from experimental collaborations like MiniBooNE Aguilar-Arevalo and others (2018), LSND Auerbach and others (2001); Batell et al. (2009), E137 Batell et al. (2014b), BaBar Lees and others (2017), and NA64 Andreev and others (2023), currently driving the leading bound on this scenario. Note that phenomenological limits from KARMEN data might cover some of the regions covered by our analyses Dutta et al. (2023). All these limits, however, should be interpreted with care, as they probe different processes and interaction channels. In particular, they are not directly related to CEνNS. MiniBooNE and LSND, although neutrino-beam experiments, are based on quasi-elastic scattering and inverse $\beta$ decay, respectively, while E137 and BaBar rely on $e^{-}-\text{Al}$ fixed-target and $e^{+}e^{-}$ collider searches. Consequently, the comparison is indicative rather than strictly equivalent. We further show the preferred curves in the $Y$ parameter corresponding to the observed relic DM abundance for scalar, pseudo-Dirac, and Majorana dark matter Berlin et al. (2019). As discussed above, these lines represent well-motivated target sensitivities: below them, the DM annihilation cross section is insufficient to reduce the relic abundance to the observed value. Above them, the DM would be underabundant and require an additional candidate to explain the total observed relic density.
@@
As can be seen, future spallation-based neutrino experiments have the potential to improve upon existing constraints, particularly in the DM mass range $6~\mathrm{MeV}\lesssim m_{\phi}\lesssim 40~\mathrm{MeV}$, provided the detector is located along the proton-beam direction. In this region, the projected sensitivities can surpass the thermal relic density target for a scalar DM candidate in the case of the CsI and Xe detectors at J-PARC, with optimal performance around $m_{\phi}\sim\mathcal{O}(25~\mathrm{MeV})$. In the same mass region, the detectors at ESS and CSNS can probe new regions of the parameter space up to the relic density target. Indeed, at the ESS and J-PARC, germanium-based detectors provide the strongest constraints, especially in the low-mass regime, reaching sensitivities down to $Y\sim 2\times 10^{-12}$. This is mostly due to the low recoil-energy threshold achievable in such detectors. In the case of the CsI detector at these facilities, it is clear that doubling its volume (as expected at J-PARC Collar and others (2025)), leads to a non-negligible improvement in sensitivity reach, as visible by comparing the J-PARC vs the ESS panels in Fig. 8. Despite its much larger mass, the CsI detector at the CSNS is slightly less constraining due to its higher operation threshold together with the higher SSB background, but still it allows to probe $Y\sim 2\times 10^{-11}$ at tens of MeV dark photon masses, beyond the relic density benchmark. In the low-mass regime, sensitivities tend to saturate mostly due to detector limitations, namely the loss of efficiency at low recoil energies.
@@
On the other side of the parameter space instead, the sensitivity of spallation neutron source experiments rapidly reduces for mediator masses above $m_{A^{\prime}}\sim 100~\mathrm{MeV}$ (i.e., $m_{\phi}\gtrsim 45~\mathrm{MeV}$), reflecting the kinematic constraint for DM production through neutral pion decays. Access to heavier DM masses would require higher energy proton beams capable of producing heavier neutral mesons. Depending on the proton beam collision energy, heavier mesons can be produced, thereby extending the experimental reach to larger values of $m_{\phi}$. In addition, DM production via proton Bremsstrahlung could further expand the accessible parameter space deNiverville et al. (2017); Blümlein and Brunner (2014). Extending up to the energy regime of SHIP and FASER, the Drell-Yan production channel is also available. Spallation neutron source facilities provide a unique production channel that facilitates the “deconstruction” of the physics underlying a DM signal. High statistics combined with high background discrimination power, enabled by timing spectra, make these facilities a unique environment for light secluded sectors. Altogether, these considerations highlight the strengths of spallation neutron source facilities in probing light dark sectors. Information stemming from these measurements are therefore complementary to those at high-energy accelerator facilities.
@@
Next, in Fig. 9 we compare our projected exclusion reach in the $(m_{\phi},Y)$ plane for the dark photon portal scenario with the sensitivities expected at other future facilities. Our projections assume on-axis detector configurations and three years of data taking. We include sensitivity estimates for two proposed upgrades of COHERENT detectors optimized for CEνNS detection: a 700 kg undoped cryogenic CsI detector Barbeau and others (2024) (pink, dashed) and a 10-tonne LAr detector Barbeau and others (2024) (pink, dotted). We also show projections for the first phase of LDMX Åkesson and others (2018); Batell and others (2022) (light blue, dashed), the movable near detector of DUNE-PRISM with a 75-tonne LAr target De Romeri et al. (2019) (light green, dashed), and a five-year run of SHiP Ahdida and others (2021); Batell and others (2022) (coral, dashed). These comparisons should also be interpreted with care. LDMX and SHiP are dedicated DM experiments, while DUNE operates at higher beam energies and primarily probes $\text{DM}-e^{-}$ scattering. Moreover, the projected sensitivities of the future COHERENT upgrades benefit from substantially larger detector masses. Despite this, our results show that experiments at ESS, J-PARC, and CSNS can competitively explore a broad region of parameter space. In particular, they are sensitive to DM masses in the range $m_{\phi}\simeq 20$4040 MeV, touching or surpassing the thermal relic target for scalar DM.
@@
All the sensitivities presented up to this point in this section have been derived from the DM event rate computed using our GEANT4-based simulations, as described in Sec. IV. As anticipated, for the ESS-Xe detector we repeated the analysis employing instead the neutral pion PDF obtained from the S & W parametrization. Although not shown here, we have verified that their spectral shapes are remarkably consistent between the two approaches, with their normalization differing by less than a factor of 1.3. Overall, the S & W parametrization systematically predicts fewer events than GEANT4, leading to slightly more conservative sensitivity projections. We present the result of this comparison in Fig. 10 where we compare the projected sensitivities for the Xe detector at the ESS obtained using the two approaches discussed in this work: the neutral pion PDF derived from the S & W parametrization and that extracted from our dedicated GEANT4 simulations. The overall agreement between the two approaches propagates to the final exclusion contours, which are found to be notably similar over all the mass range probed with our analysis. The differences between the two approaches amount to less than 15% in the final sensitivities and are barely visible in the resulting plots. It is worth noting that larger uncertainties may arise from the simulation framework itself, for instance, when switching between hadronic models in GEANT4.
@@ VI Conclusions
We have considered representative detector configurations, including planned gaseous xenon (20 kg), high-purity germanium (7 kg), and cryogenic CsI scintillating detectors (22.5 kg and 44 kg) at ESS and J-PARC, as well as a 300 kg CsI detector under construction at CSNS. A detailed discussion of the neutral pion flux calculation at these facilities has been presented. We compared two approaches: a dedicated GEANT4 simulation of proton collisions on mercury and tungsten targets, and the Sanford-Wang parametrization of charged-pion differential cross sections. We find that both methods yield consistent results within the intrinsic uncertainties of the overall calculation. Our analysis focuses on a sub-GeV scalar dark matter candidate in the minimal dark photon portal framework, as well as on a baryophilic scenario based on gauged baryon number. Since the considered detectors are optimized for nuclear recoil measurements, we have studied coherent elastic dark matter-nucleus scattering signals.
Projected sensitivities have been presented in the $m_{\phi}$-$Y$ plane. Our results show that spallation-source neutrino experiments can probe interesting regions of parameter space, particularly in the DM mass range $10~\mathrm{MeV}\lesssim m_{\phi}\lesssim 40~\mathrm{MeV}$. Sensitivities reach or extend beyond the scalar DM thermal relic target . Even stronger sensitivities are obtained in baryophilic mediator scenarios (gauged baryon number models), as detailed in App. A. Our results highlight the strong complementarity between spallation-source facilities, optimized for CEνNS measurements, and other accelerator-based dark matter searches. By leveraging their nuclear recoil capabilities, these experiments offer a uniquely powerful environment for dark matter searches in the MeV mass range.